OpenStudy (anonymous):

Write an expression that could be use to change 5 revolution per minute to radian per second.

OpenStudy (amistre64):

\[\frac{5rev}{1min}\frac{1min}{60sec}\frac{2pi\ rads}{1rev}\]

OpenStudy (amistre64):

1min = 60sec 1min/60sec is an equivalent statement that means "1"; when we multiply by "1" the value never changes, only the way it looks

OpenStudy (amistre64):

2pi rads = 1 revolutions; 2pi rads/1rev is a useful form of "1"

OpenStudy (amistre64):

cancel "like terms" and simplify

OpenStudy (anonymous):

than what answer?

OpenStudy (amistre64):

"Write an expression that could be use to change" that IS the answer; the just change 5 revs to N revs and 1min to Mmin and cancel out "like terms"

OpenStudy (anonymous):

2pi rads = 1 revolutions than 5 revolutions = 10 pi radi = 1800?

OpenStudy (amistre64):

\[\frac{5\cancel{rev}}{1\cancel{min}}\frac{1\cancel{min}}{60sec}\frac{2pi\ rads}{1\cancel{rev}}\] \[\frac{5(2pi)\ rads}{60\ sec}\] \[\frac{(2pi)\ rads}{60\ sec}X\]

OpenStudy (amistre64):

10pi rads in 1 minute then 10/60 = pi/6 rads per sec

OpenStudy (anonymous):

6 radian=6(180)=1080?

OpenStudy (amistre64):

pi radians = 180 degrees \[\frac{pi\ radians}{180\ degrees}= 1\] \[\frac{180\ degrees}{pi\ radians}= 1\] \[\frac{6\ radians}{1}*\frac{180\ degrees}{pi\ radians}\] \[\frac{6\ \cancel{radians}}{1}*\frac{180\ degrees}{pi\ \cancel{radians}}\] \[\frac{6*180}{pi}\ degrees\]

OpenStudy (amistre64):

you have to divide out the "pi" that is inherent in the radian measure; so it isnt just 1080; its 1080/pi

OpenStudy (anonymous):

amistre please! come back I need help

OpenStudy (amistre64):

you can always type up a new question for others to engage in as well :)

OpenStudy (anonymous):

I don't get how you caculator from 5(2pi ) ra/60 second ?

OpenStudy (amistre64):

its all about changing the way it looks; your redressing the problem, and not changing the value

OpenStudy (amistre64):

we redress it by substituting in equivalent expressions ....

OpenStudy (anonymous):

sorry I don't get how you come up 1080/pi

OpenStudy (anonymous):

I'm slow

OpenStudy (amistre64):

to change from radians to degrees; we have to divide out a pi; radians always include a pi, whether you see it or not.

OpenStudy (amistre64):

6 radians has a "pi" built into it simply because it is a radian measure

OpenStudy (anonymous):

5radi=1800? 2pi=360 1800(360)/60 =1080?

OpenStudy (amistre64):

lets start by stating the problem; that way i can stay on track in my head :) Write an expression that could be use to change 5 revolution per minute to radian per second.

OpenStudy (amistre64):

5 turns around means 5(2pi)radians = 10pi radians, right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

10pi radians every 1 minute right?

OpenStudy (amistre64):

1 minute = 60 seconds; so 1min/60secs = 1; agree?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

(10pi rads every 1 minute) ; so if we divide this by 60 seconds we get: 10pi rads 1 minute 10pi rads -------- every ------- = --------- every time, right? 60 sec 60 sec 60 sec

OpenStudy (anonymous):

than 10pi/60 second?

OpenStudy (amistre64):

reduce it down to 1 sec; 10pi rads 1pi rad -------- = ------- = (pi/6 rads) every second 60 sec 6 sec

OpenStudy (anonymous):

I don't get last part

OpenStudy (amistre64):

do you agree that we can get it down to: pi rads every 6 seconds at least?

OpenStudy (amistre64):

5(2pi rads) every 1 minute 10pi rads every 1 minute; 1 minute = 60 seconds 10pi rads every 60 seconds ; reduce pi rads every 6 seconds .... you agree?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

good; but the question ask us for every 1 second; not every 6 seconds right?; so we have to reduce again by dividing at all by 6 5(2pi rads) every 1 minute 10pi rads every 1 minute; 1 minute = 60 seconds 10pi rads every 60 seconds ; reduce pi rads every 6 seconds --- --- 6 6 pi/6 rads every 1 second

OpenStudy (anonymous):

answer=6?

OpenStudy (amistre64):

\(\cfrac{pi}{6}\) rads every 1 second, would be the answer; not "6"

OpenStudy (anonymous):

wait =30

OpenStudy (anonymous):

30?

OpenStudy (amistre64):

..... what are you seeing that makes you think the answer is anything other than \(\cfrac{pi}{6}\) ?

OpenStudy (anonymous):

pi=180? 180/6=30

OpenStudy (amistre64):

pi does not equal 180; pi = 3.14159...

OpenStudy (amistre64):

ahhh ..... then yes

OpenStudy (amistre64):

pi/6 = 30 degrees :)

OpenStudy (amistre64):

\[\frac{pi}{6}radians=\frac{180}{6}degrees=30degrees\]

OpenStudy (anonymous):

I'm confuse pi=180 pi=3.14 any way thank you fo be patient

OpenStudy (amistre64):

recall when we did: 1 minute = 60 seconds? that is a relationship between minutes and seconds. 1m = 60s is true if we use the relationship between minutes and seconds; the same thing holds for the relationship between radians and degrees; (pi)r = 180d only means something when we are using it in this context.

OpenStudy (amistre64):

if we ignore the relationship and strip away the equivalency; it becomes nonesense: 1 = 60 is nonsense on its own pi = 180 is nonsense on it own

OpenStudy (anonymous):

ty, I get it now